# Variational Search Distributions

### v2

@@ B.1 Variational distributions
where $x_{m}\in\mathcal{V}$ and $\phi_{m}\in^{|\mathcal{V}|}$. However, this simple mean-field distribution was not capable of generating convincing handwritten digits. We have also tested a variety of transition variational distributions,
@@
where $\textrm{NN}_{m}(\mathbf{x}_{t-1},\phi)$ is the $m^{\textrm{th}}$ vector output of a neural network that takes a sequence from the previous round, $\mathbf{x}_{t-1}$, as input. We have implemented multiple neural net encoder/decoder architectures for $\textrm{NN}_{m}(\mathbf{x}_{t-1},\phi)$, but we did not consider architectures of the form $\textrm{NN}_{m}(\phi)$ since the variational distribution in Equation 20 can always learn a $\phi_{m}=\textrm{NN}_{m}(\phi^{\prime})$. We found that none of these transition architectures significantly outperformed the mean-field distribution (Equation 20) when it was initialized well (e.g. fit to the CPE training sequences), see Sec. C.3 for results. We also implemented auto-regressive variational distributions of the form,
@@
For a LSTM RNN and a decoder-only transformer with a causal mask, for the latter see Phuong & Hutter (2022, Algorithm 10 & Algorithm 14) for maximum likelihood training and sampling implementation details respectively. For the digits experiment, the LSTM uses 5 layers of dimension 128, and the transformer uses 8-attention heads, with 4 layers and a feed-forward network size of 256. For the AAV and GFP experiments, the LSTM uses 4 layers of dimension 32, and the transformer uses 4 attention heads for GFP, 2 attention heads for AAV, 1 layer, and a feed-forward network size of 64. Any larger than this, and we found these models would over-fit. We use additive positional encoding for both of these models.
B.2 Class Probability Estimator Architectures
For the fitness landscape experiments on the smaller combinatorially complete datasets we use a two-hidden layer MLP, with an input embedding layer. The architecture is given in Figure 5 (a). For the larger dimensional AAV and GFP datasets, we use the convolutional architecture given in Figure 5 (b). Five fold cross validation is used to select the hyper parameters before the CPE s are trained on the whole training set for use in the subsequent experimental rounds. Model updates are performed by retraining on the whole query set.
@@ B.2 Class Probability Estimator Architectures
Figure 5: CPE architectures used for the experiments in PyTorch syntax. $\texttt{A}=|\mathcal{V}|$, $\texttt{M}=M$, GFP uses a max pooling kernel size and stride of 4, all other datasets use 2.
Appendix C Additional Experimental Results
C.1 Fitness Landscapes – Gaussian Process Probability of Improvement
Here we present additional fitness landscape experimental results, where we have used a GP as a surrogate model for $p({y}|{\mathbf{x},\mathcal{D}_{N}})$ in conjunction with a complementary Normal CDF as the PI acquisition function. This is the one of the main frameworks supported by our theoretical analysis. VSD, DbAS, CbAS and BORE make use of the GP-PI acquisition function, and so BORE is BOPR in this instance since we are not using a CPE. PEX and AdaLead only use the GP surrogate, as per their original formulation. The GP uses a simple categorical kernel with automatic relevance determination from Balandat et al. (2020),
$k(\mathbf{x},\mathbf{x}^{\prime})=\sigma\exp\!\left(-\frac{1}{M}\sum^{M}_{m=1}\frac{\mathbbm{1}[{x_{m}=x^{\prime}_{m}}]}{l_{m}}\right),$ (23)
where $\sigma$ and $l_{m}$ are hyper-parameters controlling scale and length-scale respectively. See Figure 6 for the results.
@@ C.1 Fitness Landscapes – Gaussian Process Probability of Improvement
Figure 6: Fitness landscape results using GP-PI. Precision (Equation 15), recall (Equation 16) and performance (Equation 17) – higher is better – for the combinatorially (near) complete datasets, DHFR and TrpB and TFBIND8. The random method is implemented by drawing $B$ samples uniformly.
C.2 Diversity Scores
The diversity of batches of candidates is a common thing to report in the literature, and to that end we present the diversity of our results here. We have taken the definition of diversity from (Jain et al., 2022) as,
@@ C.2 Diversity Scores
where $\mathrm{Lev}:\mathcal{X}\times\mathcal{X}\to\mathbb{N}_{0}$ is the Levenshtein distance. We caution the reader as to the interpretation of these results however, as more diverse batches often do not lead to better performance, precision, recall or simple regret (as can be seen from the Random method results). Though insufficient diversity can also explain poor performance, as in the case of BORE. Results for the fitness landscape experiment are presented in Figure 7, and black-box optimization in Figure 8.
@@
Figure 7: Fitness landscape diversity results. Higher is more diverse, as defined by Equation 24.
@@
Figure 8: Black-box optimization results for diversity on GFP and AAV with independent and auto-regressive variational distributions. Higher is more diverse, as defined by Equation 24. The PEX and AdaLead results are replicated between the plots, since they are unaffected by choice of variational distribution.
C.3 Ablations – Variational and Prior Distributions
In Figure 9 we present ablation results for VSD using different priors and variational distributions. We use the BBO experimental datasets for this task as they are higher-dimensional and so more sensitive to these design choices. We test the following prior and variational posterior distributions:
@@ C.3 Ablations – Variational and Prior Distributions
Figure 9: Ablation results for the AAV and GFP BBO experiments. VSD is trialed with different prior and variational posterior combinations, “I” indicates a simple independent informed prior and posterior, “IU” is the same but with a uniform prior, “LSTM” and “DTFM” are the LSTM and decoder only transformer prior and posteriors, “TCNN” and “TAE” are transition convolutional encoder-decoder and auto-encoder posteriors, with informed independent priors. See text for details.
From Figure 9 we can see that while using an uninformative prior works in the lower-dimensional fitness landscape experiments, using an informative prior is crucial for these higher dimensional problems. We found a similar result when using this uninformative prior with CbAS, or using a uniform initialization with DbAS and BORE. The methods are not able to make any significant progress within the experimental budget given. The independent and transition variational distributions achieve similar performance, whereas the auto-regressive models generally outperform all others. This is because of the LSTM and transformer’s superior generalization performance when generating sequences – measured both when training the priors (on held-out sequences) and during VSD adaptation.
Appendix D Theoretical analysis for GP-based CPEs
In this section, we present theoretical results concerning VSD and its estimates when equipped with Gaussian process regression models (Rasmussen & Williams, 2006). We show that VSD sampling distributions converge to a target distribution that characterizes the level set given by $\tau$. The approximation error mainly depends on the predictive uncertainty of the probabilistic model with respect to the true underlying function ${f\!\centerdot}$. For the analysis, we will assume that ${f\!\centerdot}$ is drawn from a Gaussian process, i.e., ${f\!\centerdot}\sim\mathcal{GP}(0,k)$, with a positive-semidefinite covariance (or kernel) function $k:\mathcal{X}\times\mathcal{X}\to$. In this case, we can show that the predictive uncertainty of the model converges (in probability) to zero as the number of observations grows. From this result, we prove asymptotic convergence guarantees for VSD equipped with GP-PI-based CPEs. These results form the basis for our analysis of CPEs based on neural networks (Appendix E).
D.1 Gaussian process posterior
@@ D.2 Background
$\forall\varepsilon>0,\quad\exists C_{\varepsilon}\in(0,\infty):\quad\mathbb{P}\!\left[|\rho_{t}|>C_{\varepsilon}\right]\leq\varepsilon,\quad\forall t\geq T_{\varepsilon}\,,$ (33)
for some $T_{\varepsilon}\in\mathbb{N}$. For almost sure convergence, we may also say that a sequence of random variables $\rho_{t}$, $t\in\mathbb{N}$, is almost surely ${\mathcal{{O}}}(g(t))$ if $\mathbb{P}\!\left[\rho_{t}\in{\mathcal{{O}}}(g(t))\right]=1$. A deeper overview on these notations and their properties can be found in García-Portugués (2024).
D.3 Auxiliary results
@@
Lemma D.1 (Chowdhury & Gopalan (2017, Appendix F)).
@@
Lemma D.2 (Second Borel-Cantelli lemma).
Let $\{{\mathcal{{A}}}_{t}\}_{t=1}^{\infty}$ be a sequence of events where ${\mathcal{{A}}}_{t}\in{\mathfrak{{F}}}_{t}$, for all $t\in\mathbb{N}$, and let $\chi_{t}:\omega\mapsto\mathbbm{1}[{\omega\in{\mathcal{{A}}}_{t}}]$, for $\omega\in\Omega$. Then the following holds with probability 1:
@@
Now assume that, for a given $t>0$, it holds that $\sigma_{t}^{2}(\mathbf{x})\leq\tilde{\sigma}_{t}^{2}(\mathbf{x})$, for all $\mathbf{x}\in\mathcal{X}$. We will then check if the inequality remains valid at $t+1$. By D.1, we have that:
@@
For any $\mathbf{x}\in\mathcal{X}$ such that $\mathbf{x}\neq\mathbf{x}_{t+1}$, we know that $\tilde{k}_{t}(\mathbf{x},\mathbf{x}_{t+1})\geq 0$, so that (again by D.1):
@@
$\forall t\geq 1,\quad\frac{1}{\tilde{\sigma}_{t}^{2}(\mathbf{x})}=\frac{1}{\tilde{\sigma}_{t-1}^{2}(\mathbf{x}_{t})}+\frac{\mathbbm{1}[{\mathbf{x}_{t}=\mathbf{x}}]}{\sigma_{\epsilon}^{2}},\quad\forall\mathbf{x}\in\mathcal{X}.$ (48)
@@
$\forall t\geq 1,\quad\frac{1}{\tilde{\sigma}_{t}^{2}(\mathbf{x})}=\frac{1}{\tilde{\sigma}_{0}^{2}(\mathbf{x})}+\frac{1}{\sigma_{\epsilon}^{2}}\sum_{i=1}^{t}\mathbbm{1}[{\mathbf{x}_{i}=\mathbf{x}}]\,,\quad\forall\mathbf{x}\in\mathcal{X}\,.$ (49)
The result in D.3 then follows as the reciprocal of the above, which concludes the proof. ∎
Lemma D.4.
@@ D.4 Asymptotic convergence
for a some sequence of lower bounds $\{B_{t}\}_{t\in\mathbb{N}}$. Then, under D.1, given observations at $\{\mathbf{x}_{i}\}_{i=1}^{t}$, the following holds with probability 1:
@@
At any iteration $t$, the posterior variance $\sigma_{t}^{2}$ of a GP model is upper bounded by a worst case assumption of no correlation between observations (see D.3). In this case, we have that:
@@
The only random variable to be bounded in Equation 54 is $N_{t}$. Let $\chi_{t}:=\mathbbm{1}[{\mathbf{x}_{t}=\mathbf{x}}]$, so that:
$N_{t}=\sum_{i=1}^{t}\chi_{i}=\sum_{i=1}^{t}\mathbbm{1}[{\mathbf{x}_{t}=\mathbf{x}}]\,,\quad t\geq 1.$ (55)
We now apply the second Borel-Cantelli lemma (D.2) to $N_{t}$. Namely, let $\widehat{N}_{t}$ denote the sum of conditional expectations of $\{\chi_{i}\}_{i=1}^{t}$ given available data, i.e.:
$\widehat{N}_{t}:=\sum_{i=1}^{t}\mathbb{E}[\chi_{i}\mid{\mathfrak{{F}}}_{i-1}]=\sum_{i=1}^{t}\mathbb{E}[\mathbbm{1}[{\mathbf{x}_{t}=\mathbf{x}}]\mid{\mathfrak{{F}}}_{i-1}]=\sum_{i=1}^{t}\mathbb{P}\!\left[\mathbf{x}_{i}=\mathbf{x}\mid{\mathfrak{{F}}}_{i-1}\right]\,.$ (56)
By D.2, we know that the following holds for some $L\in$:
@@
Theorem D.1.
Let assumptions D.1 to D.4 hold. Then the following holds with probability 1 for VSD equipped with GP-PI:
@@
Let $\ell_{t}(\mathbf{x}):=p({y>\tau_{t}}|{\mathbf{x},\mathcal{D}_{t}})$. For any given $\mathbf{x}\in\mathcal{X}$ where $p({\mathbf{x}})>0$, by D.2, we have that the next candidate will be sampled according to:
@@
where we used the fact that $\mathbb{E}_{p({\mathbf{x}})}\!\left[{\ell_{t}(\mathbf{x})}\right]\leq 1$, since $\ell_{t}(\mathbf{x})\leq 1$, for all $\mathbf{x}\in\mathcal{X}$. As $p({\mathbf{x}})>0$, we only have to derive a lower bound on $\ell_{t}(\mathbf{x})$ to apply D.5 and derive a convergence rate.
@@
where $\Psi(\cdot)$ denotes the cumulative distribution function of a standard normal random variable, and $\lVert\cdot\rVert_{\infty}$ denotes the essential supremum of a function under $\mathbb{P}$ (the probability measure of the underlying abstract probability space). Therefore, if $\lim_{t\to\infty}\lVert\mu_{t}\rVert_{\infty}<\infty$, we will have that $\lim_{t\to\infty}\ell_{t}(\mathbf{x})>0$, and the sum in D.5 will diverge.
@@
As $\mathbb{E}[\mathbb{E}[\lVert{f\!\centerdot}\rVert_{\infty}\mid{\mathfrak{{F}}}_{t}]]=\mathbb{E}[\lVert{f\!\centerdot}\rVert_{\infty}]<\infty$ (cf. D.4), an application of Markov's inequality implies that:
@@
Hence, for any $\varepsilon_{a}\in(0,b_{\infty}(a))$, there is $N_{a}\in\mathbb{N}$, such that $\mathbb{P}\!\left[\mathbf{x}_{t}=\mathbf{x}\mid{\mathfrak{{F}}}_{t-1}\right]\geq b_{\infty}(a)-\varepsilon_{a}>0$, for all $t\geq N_{a}$. As a result, $\sum_{t^{\prime}=1}^{t}\mathbb{P}\!\left[\mathbf{x}_{t^{\prime}}=\mathbf{x}\mid{\mathfrak{{F}}}_{t^{\prime}-1}\right]\geq(b_{\infty}(a)-\varepsilon_{a})(t-N_{a})$, for all $t\geq N_{a}$, which asymptotically diverges at a rate proportional to $t$. By D.5 and the definition of the big-${\mathcal{{O}}}$ notation, for any $\mathbf{x}\in\mathcal{X}$, we then have that:
@@
Remark D.1.
The convergence rate in Theorem D.1 is optimal and cannot be further improved. As shown by previous works in the online learning literature (Mutný & Krause, 2018; Takeno et al., 2024), a lower bound on the GP variance at each iteration $t\geq 1$ is given by $\sigma_{t}^{2}(\mathbf{x})\geq\sigma_{\epsilon}^{2}(\sigma_{\epsilon}^{2}+t)^{-1}$ (assuming $k(\mathbf{x},\mathbf{x})=1$), which is the case when every observation in the dataset was collected at the same point $\mathbf{x}\in\mathcal{X}$ (see Takeno et al., 2024, Lem. 4.2). Therefore, the lower and upper bounds on the asymptotic convergence rates for the GP variance differ by only up to a multiplicative constant.
The result in Theorem D.1 now allows us to derive a convergence rate for VSD's approximations to the level-set distributions. To do so, however, we will require the following mild assumption, which is satisfied by any prior distribution which has support on the entire domain $\mathcal{X}$.
Assumption D.5.
@@
We first prove an upper bound for the KL divergence in terms of the PI approximation error. We then derive a bound for this term and apply Theorem D.1 to obtain a convergence rate.
@@
By D.4, we know that $\mathbb{E}[\lVert{f\!\centerdot}\rVert_{\infty}]<\infty$, which implies that $\mathbb{P}\!\left[\lVert{f\!\centerdot}\rVert_{\infty}<\infty\right]=1$ by Markov's inequality. Next, we derive a bound for the approximation error term.
@@
Applying Theorem D.1, we then have that:

### v4

@@ C.3 Variational Distributions
where $x_{m}\in\mathcal{V}$ and $\phi_{m}\in^{|\mathcal{V}|}$. However, this simple mean-field distribution was not capable of generating convincing handwritten digits or, in some cases, higher-dimensional sequences. We have also tested a variety of transition variational distributions,
@@
where $\textrm{NN}_{m}(\mathbf{x}_{t-1},\phi)$ is the $m^{\textrm{th}}$ vector output of a neural network that takes a sequence from the previous round, $\mathbf{x}_{t-1}$, as input. We have implemented multiple neural net encoder/decoder architectures for $\textrm{NN}_{m}(\mathbf{x}_{t-1},\phi)$, but we did not consider architectures of the form $\textrm{NN}_{m}(\phi)$ since the variational distribution in Equation 20 can always learn a $\phi_{m}=\textrm{NN}_{m}(\phi^{\prime})$. We found that none of these transition architectures significantly outperformed the mean-field distribution (Equation 20) when it was initialized well (e.g. fit to the CPE training sequences), see Sec. D.4 for results. We also implemented auto-regressive variational distributions of the form,
@@
For a LSTM RNN and a decoder-only transformer with a causal mask, for the latter see Phuong & Hutter (2022, Algorithm 10 & Algorithm 14) for maximum likelihood training and sampling implementation details respectively. We list the configurations of the LSTM and transformer variational distributions in Table C.2. We use additive positional encoding for all of these models. When using these models for priors or initialization of variational distributions, we find that over-fitting can be an issue. To circumvent this, we use early stopping for larger training datasets, or data augmentation techniques for smaller training datasets (as in the case of the Ehrlich functions).
Configuration Digits AAV GFP Ehrlich 15 Ehrlich 32 Ehrlich 64
LSTM Layers 5 4 4 3 3 3
Network size 128 32 32 32 32 64
Embedding size 4 10 10 10 10 10
Transformer Layers 4 1 1 2 2 2
Network Size 256 64 64 32 64 128
Attention heads 8 2 2 1 2 3
Embedding size 32 20 20 10 20 30
Table C.2: LSTM and transformer network configuration.
C.4 Class Probability Estimator Architectures
For the fitness landscape experiments on the smaller combinatorially complete datasets we use a two-hidden layer MLP, with an input embedding layer. The architecture is given in Figure C.2 (a). For the larger dimensional AAV and GFP datasets and Ehrlich function benchmark, we use the convolutional architecture given in Figure C.2 (b). On all but the Ehrlich benchmark, five fold cross validation was used to select the hyper parameters before the CPE s are trained on the whole training set for use in the subsequent experimental rounds. For the Ehrlich benchmark we do not use cross-validation to select the CPE hyper parameters – but we do use an additive ensemble of 10 randomly initialized CNNs for the CPE following LaMBO-2. Model updates are performed by retraining on the whole query set.
@@ C.4 Class Probability Estimator Architectures
Figure C.2: CPE architectures used for the experiments in PyTorch syntax. $\texttt{A}=|\mathcal{V}|$, $\texttt{M}=M$, GFP uses a max pooling kernel size and stride of 4, all other datasets and benchmarks use 2. The Ehrlich function benchmark uses and ensemble of 10 randomly initialized CNNs that are additively combined. The Ehrlich-15 functions use a kernel size of 3, all other BBO experiments use a kernel size of 7. LaMBO-2 uses the same kernel size as our CNNs for the Ehrlich functions.
Appendix D Additional Experimental Results
D.1 Fitness Landscapes – Gaussian Process Probability of Improvement
Here we present additional fitness landscape experimental results, where we have used a GP as a surrogate model for $p({y}|{\mathbf{x},\mathcal{D}_{N}})$ in conjunction with a complementary Normal CDF as the PI acquisition function. This is one of the main frameworks supported by our theoretical analysis. VSD, DbAS, CbAS and BORE can make use of the GP-PI acquisition function, and so BORE is BOPR (Daulton et al., 2022) in this instance since we are not using a CPE. PEX and AdaLead only use the GP surrogate, as per their original formulation. The GP uses a simple categorical kernel with automatic relevance determination from Balandat et al. (2020),
$k(\mathbf{x},\mathbf{x}^{\prime})=\sigma\exp\!\left(-\frac{1}{M}\sum^{M}_{m=1}\frac{\mathds{1}[{x_{m}=x^{\prime}_{m}}]}{l_{m}}\right),$ (23)
where $\sigma$ and $l_{m}$ are hyper-parameters controlling scale and length-scale respectively. See Figure D.3 for the results.
@@ D.1 Fitness Landscapes – Gaussian Process Probability of Improvement
Figure D.3: Fitness landscape results using GP-PI. Precision (Equation 15), recall (Equation 16) and performance (Equation 17) – higher is better – for the combinatorially (near) complete datasets, DHFR and TrpB and TFBIND8. The random method is implemented by drawing $B$ samples uniformly.
D.2 Ehrlich Function HOLO Results
See Figure D.4 for BBO results on the original holo Ehrlich function implementation (Stanton et al., 2024). We present additional diversity scores for these and the poli implementation in Sec. D.3.
(a) $M=15$
(b) $M=32$
(c) $M=64$
Figure D.4: Ehrlich function (holo implementation) BBO results. VSD and CbAS with different variational distributions; mean field (MF), LSTM and transformer (TFM), compared against genetic algorithm (GA) and LaMBO-2 baselines.
D.3 Diversity Scores
The diversity of batches of candidates is a common thing to report in the literature, and to that end we present the diversity of our results here. We have taken the definition of pair-wise diversity from (Jain et al., 2022) as,
@@ D.3 Diversity Scores
where $\mathrm{Lev}:\mathcal{X}\times\mathcal{X}\to\mathbb{N}_{0}$ is the Levenshtein distance. We caution the reader as to the interpretation of these results however, as more diverse batches often do not lead to better performance, precision, recall or simple regret (as can be seen from the Random method results). Though insufficient diversity can also explain poor performance, as in the case of BORE. Results for the fitness landscape experiment are presented in Figure D.5, and black-box optimization for AAV & GFP in Figure D.6 and Ehrlich functions in Figure D.7.
@@
Figure D.5: Fitness landscape diversity results. Higher is more diverse, as defined by Equation 24.
@@
Figure D.6: Black-box optimization results for diversity on GFP and AAV with independent and auto-regressive variational distributions. Higher is more diverse, as defined by Equation 24. The PEX and AdaLead results are replicated between the plots, since they are unaffected by choice of variational distribution.
poli
holo
(a) $M=15$
(b) $M=32$
(c) $M=64$
Figure D.7: Black-box optimization results for diversity on the poli and holo implementations of the Ehrlich functions. Higher is more diverse, as defined by Equation 24.
D.4 Ablations – Variational and Prior Distributions
In Figure D.8 we present ablation results for VSD using different priors and variational distributions. We use the BBO experimental datasets for this task as they are higher-dimensional and so more sensitive to these design choices. We test the following prior and variational posterior distributions:
@@ D.4 Ablations – Variational and Prior Distributions
Figure D.8: Ablation results for the AAV and GFP BBO experiments. VSD is trialed with different prior and variational posterior combinations, “I” indicates a simple independent informed prior and posterior, “IU” is the same but with a uniform prior, “LSTM” and “DTFM” are the LSTM and decoder only transformer prior and posteriors, “TCNN” and “TAE” are transition convolutional encoder-decoder and auto-encoder posteriors, with informed independent priors. See text for details.
From Figure D.8 we can see that while using an uninformative prior works in the lower-dimensional fitness landscape experiments, using an informative prior is crucial for these higher dimensional problems. We found a similar result when using this uninformative prior with CbAS, or using a uniform initialization with DbAS and BORE. The methods are not able to make any significant progress within the experimental budget given. The independent and transition variational distributions achieve similar performance, whereas the auto-regressive models generally outperform all others. This is because of the LSTM and transformer’s superior generalization performance when generating sequences – measured both when training the priors (on held-out sequences) and during VSD adaptation.
Appendix E Theoretical analysis for GP-based CPEs
In this section, we present theoretical results concerning VSD and its estimates when equipped with Gaussian process regression models (Rasmussen & Williams, 2006). We show that VSD sampling distributions converge to a target distribution that characterizes the level set given by $\tau$. The approximation error mainly depends on the predictive uncertainty of the probabilistic model with respect to the true underlying function ${f\!\centerdot}$. For the analysis, we will assume that ${f\!\centerdot}$ is drawn from a Gaussian process, i.e., ${f\!\centerdot}\sim\mathcal{GP}(0,k)$, with a positive-semidefinite covariance (or kernel) function $k:\mathcal{X}\times\mathcal{X}\to$. In this case, we can show that the predictive uncertainty of the model converges (in probability) to zero as the number of observations grows. From this result, we prove asymptotic convergence guarantees for VSD equipped with GP-PI-based CPEs. These results form the basis for our analysis of CPEs based on neural networks (Appendix F).
E.1 Gaussian process posterior
@@ E.2 Background
$\forall\varepsilon>0,\quad\exists C_{\varepsilon}\in(0,\infty):\quad\mathbb{P}\!\left[\frac{|\rho_{t}|}{g(t)}>C_{\varepsilon}\right]\leq\varepsilon,\quad\forall t\geq T_{\varepsilon}\,,$ (33)
for some $T_{\varepsilon}\in\mathbb{N}$. For almost sure convergence, we may also say that a sequence of random variables $\rho_{t}$, $t\in\mathbb{N}$, is almost surely ${\mathcal{{O}}}(g(t))$ if $\mathbb{P}\!\left[\rho_{t}\in{\mathcal{{O}}}(g(t))\right]=1$. A deeper overview on these notations and their properties can be found in García-Portugués (2024). In addition, as it is common in the multi-armed bandits literature, we use the variants $\widetilde{{\mathcal{{O}}}}$ and $\widetilde{{\mathcal{{O}}}}_{\mathbb{P}}$ to denote asymptotic rates which are valid up to logarithmic factors.
E.3 Auxiliary results
@@
Lemma E.1 (Chowdhury & Gopalan (2017, Appendix F)).
@@
Lemma E.2 (Second Borel-Cantelli lemma).
Let $\{{\mathcal{{A}}}_{t}\}_{t=1}^{\infty}$ be a sequence of events where ${\mathcal{{A}}}_{t}\in{\mathfrak{{F}}}_{t}$, for all $t\in\mathbb{N}$, and let $\chi_{t}:\omega\mapsto\mathds{1}[{\omega\in{\mathcal{{A}}}_{t}}]$, for $\omega\in\Omega$. Then the following holds with probability 1:
@@
Now assume that, for a given $t>0$, it holds that $\sigma_{t}^{2}(\mathbf{x})\leq\tilde{\sigma}_{t}^{2}(\mathbf{x})$, for all $\mathbf{x}\in\mathcal{X}$. We will then check if the inequality remains valid at $t+1$. By E.1, we have that:
@@
For any $\mathbf{x}\in\mathcal{X}$ such that $\mathbf{x}\neq\mathbf{x}_{t+1}$, we know that $\tilde{k}_{t}(\mathbf{x},\mathbf{x}_{t+1})\geq 0$, so that (again by E.1):
@@
$\forall t\geq 1,\quad\frac{1}{\tilde{\sigma}_{t}^{2}(\mathbf{x})}=\frac{1}{\tilde{\sigma}_{t-1}^{2}(\mathbf{x}_{t})}+\frac{\mathds{1}[{\mathbf{x}_{t}=\mathbf{x}}]}{\sigma_{\epsilon}^{2}},\quad\forall\mathbf{x}\in\mathcal{X}.$ (48)
@@
$\forall t\geq 1,\quad\frac{1}{\tilde{\sigma}_{t}^{2}(\mathbf{x})}=\frac{1}{\tilde{\sigma}_{0}^{2}(\mathbf{x})}+\frac{1}{\sigma_{\epsilon}^{2}}\sum_{i=1}^{t}\mathds{1}[{\mathbf{x}_{i}=\mathbf{x}}]\,,\quad\forall\mathbf{x}\in\mathcal{X}\,.$ (49)
The result in E.3 then follows as the reciprocal of the above, which concludes the proof. ∎
Lemma E.4.
@@ E.4 Asymptotic convergence
for a some sequence of lower bounds $\{B_{t}\}_{t\in\mathbb{N}}$. Then, under E.1, given observations at $\{\mathbf{x}_{i}\}_{i=1}^{t}$, the following holds with probability 1:
@@
At any iteration $t$, the posterior variance $\sigma_{t}^{2}$ of a GP model is upper bounded by a worst case assumption of no correlation between observations (see E.3). In this case, we have that:
@@
The only random variable to be bounded in Equation 54 is $N_{t}$. Let $\chi_{t}:=\mathds{1}[{\mathbf{x}_{t}=\mathbf{x}}]$, so that:
$N_{t}=\sum_{i=1}^{t}\chi_{i}=\sum_{i=1}^{t}\mathds{1}[{\mathbf{x}_{t}=\mathbf{x}}]\,,\quad t\geq 1.$ (55)
We now apply the second Borel-Cantelli lemma (E.2) to $N_{t}$. Namely, let $\widehat{N}_{t}$ denote the sum of conditional expectations of $\{\chi_{i}\}_{i=1}^{t}$ given available data, i.e.:
$\widehat{N}_{t}:=\sum_{i=1}^{t}\mathbb{E}[\chi_{i}\mid{\mathfrak{{F}}}_{i-1}]=\sum_{i=1}^{t}\mathbb{E}[\mathds{1}[{\mathbf{x}_{t}=\mathbf{x}}]\mid{\mathfrak{{F}}}_{i-1}]=\sum_{i=1}^{t}\mathbb{P}\!\left[\mathbf{x}_{i}=\mathbf{x}\mid{\mathfrak{{F}}}_{i-1}\right]\,.$ (56)
By E.2, we know that the following holds for some $L\in$:
@@
Theorem E.1.
Let assumptions E.1 to E.4 hold. Then the following holds with probability 1 for VSD equipped with GP-PI:
@@
Let $\ell_{t}(\mathbf{x}):=p({y>\tau_{t}}|{\mathbf{x},\mathcal{D}_{t}})$. For any given $\mathbf{x}\in\mathcal{X}$ where $p({\mathbf{x}})>0$, by E.2, we have that the next candidate will be sampled according to:
@@
where we used the fact that $\mathbb{E}_{p({\mathbf{x}})}\!\left[{\ell_{t}(\mathbf{x})}\right]\leq 1$, since $\ell_{t}(\mathbf{x})\leq 1$, for all $\mathbf{x}\in\mathcal{X}$. As $p({\mathbf{x}})>0$, we only have to derive a lower bound on $\ell_{t}(\mathbf{x})$ to apply E.5 and derive a convergence rate.
@@
where $\Psi(\cdot)$ denotes the cumulative distribution function of a standard normal random variable, and $\lVert\cdot\rVert_{\infty}$ denotes the essential supremum of a function under $\mathbb{P}$ (the probability measure of the underlying abstract probability space). Therefore, if $\lim_{t\to\infty}\lVert\mu_{t}\rVert_{\infty}<\infty$, we will have that $\lim_{t\to\infty}\ell_{t}(\mathbf{x})>0$, and the sum in E.5 will diverge.
@@
As $\mathbb{E}[\mathbb{E}[\lVert{f\!\centerdot}\rVert_{\infty}\mid{\mathfrak{{F}}}_{t}]]=\mathbb{E}[\lVert{f\!\centerdot}\rVert_{\infty}]<\infty$ (cf. E.4), an application of Markov's inequality implies that:
@@
Hence, for any $\varepsilon_{a}\in(0,b_{\infty}(a))$, there is $N_{a}\in\mathbb{N}$, such that $\mathbb{P}\!\left[\mathbf{x}_{t}=\mathbf{x}\mid{\mathfrak{{F}}}_{t-1}\right]\geq b_{\infty}(a)-\varepsilon_{a}>0$, for all $t\geq N_{a}$. As a result, $\sum_{t^{\prime}=1}^{t}\mathbb{P}\!\left[\mathbf{x}_{t^{\prime}}=\mathbf{x}\mid{\mathfrak{{F}}}_{t^{\prime}-1}\right]\geq(b_{\infty}(a)-\varepsilon_{a})(t-N_{a})$, for all $t\geq N_{a}$, which asymptotically diverges at a rate proportional to $t$. By E.5 and the definition of the big-${\mathcal{{O}}}$ notation, for any $\mathbf{x}\in\mathcal{X}$, we then have that:
@@
Remark E.1.
The convergence rate in Theorem E.1 is optimal and cannot be further improved. As shown by previous works in the online learning literature (Mutný & Krause, 2018; Takeno et al., 2024), a lower bound on the GP variance at each iteration $t\geq 1$ is given by $\sigma_{t}^{2}(\mathbf{x})\geq\sigma_{\epsilon}^{2}(\sigma_{\epsilon}^{2}+t)^{-1}$ (assuming $k(\mathbf{x},\mathbf{x})=1$), which is the case when every observation in the dataset was collected at the same point $\mathbf{x}\in\mathcal{X}$ (see Takeno et al., 2024, Lem. 4.2). Therefore, the lower and upper bounds on the asymptotic convergence rates for the GP variance differ by only up to a multiplicative constant.
The result in Theorem E.1 now allows us to derive a convergence rate for VSD's approximations to the level-set distributions. To do so, however, we will require the following mild assumption, which is satisfied by any prior distribution which has support on the entire domain $\mathcal{X}$.
Assumption E.5.
@@
We first prove an upper bound for the KL divergence in terms of the PI approximation error. We then derive a bound for this term and apply Theorem E.1 to obtain a convergence rate.
@@
By E.4, we know that $\mathbb{E}[\lVert{f\!\centerdot}\rVert_{\infty}]<\infty$, which implies that $\mathbb{P}\!\left[\lVert{f\!\centerdot}\rVert_{\infty}<\infty\right]=1$ by Markov's inequality. Next, we derive a bound for the approximation error term.
@@
Applying Theorem E.1, we then have that: